# Bridging Earth and Space: A Survey on HAPS for Non-Terrestrial Networks

### v1

@@ III HAPS Use Cases
The unique characteristics of HAPS enable a wide range of applications across telecommunications and beyond. In telecommunications, they primarily aim to extend coverage (Section III-A) and enhance network efficiency (Section III-B). Their rapid deployment and adaptability make them essential for bridging connectivity gaps in remote, underserved, or disaster-affected regions. Serving as mobile urban BSs or backhaul providers, HAPS improve capacity, reliability, and energy efficiency. Beyond communications, they support advanced applications (Section III-C), including real-time IoT sensing, immersive augmented reality (AR)/ virtual reality (VR), precision control via URLLC (Section III-D), and energy optimization (Section III-E). By bringing processing closer to UEs, HAPS reduce latency and enable innovations across sectors such as agriculture, logistics, and disaster management. This section explores their roles in connectivity enhancement and advanced use cases shaping next-generation networks.
III-A Connectivity Enhancement
III-A1 Remote and Rural Area Coverage
@@
Collaboration between HAPS and LEO satellites was explored in [18], which formulated a mixed-integer nonlinear programming problem to enhance large-scale access and backhaul. The proposed restricted trilateral matching algorithm improved coordination, connectivity, and efficiency in remote regions.
Study [19] analyzed HAPS-LAPS integration using stochastic geometry, introducing a framework to assess platform density and altitude effects under backhaul constraints, offering design guidance for robust aerial networks in rural areas.
In [14], HAPS were studied for extending rural coverage with low UE density through integration with TNs, emphasizing coexistence via dynamic resource management, interference mitigation, and techniques such as advanced antennas, radio environment maps, and device-to-device communication.
Case study [20] optimized resource allocation for heterogeneous services in rural HAPS networks using free space optical (FSO) backhaul. It proposed static and dynamic allocation schemes leveraging DRL, improving age of information, data rates, and overall network efficiency.
III-A2 Natural Disaster Region Coverage
@@
In the aftermath of natural disasters, when TNs collapse, HAPS can provide real-time aerial surveillance and rapid, reliable communication. Covering vast and remote regions, they ensure effective disaster response and emergency connectivity for first responders [21]. Study [22] examined UAV-based HAPS for disaster scenarios, highlighting their rapid deployment, flexible coverage, and communication recovery, while identifying challenges such as backhaul limitations, platform constraints, and communication capacity that require targeted solutions. Real-time monitoring using HAPS was demonstrated in [23], showing persistent, high-resolution imaging from altitudes around 20 km, enabling stable, continuous coverage of affected areas. In [24], HAPS restored services in a disrupted universal mobile telecommunications system (UMTS) network, achieving 70% throughput recovery when terrestrial BSs were unavailable, underscoring their effectiveness in disaster recovery operations.
III-A3 Unexpected Traffic Events
@@
The inherent mobility and on-demand deployment of HAPS make them well suited for managing dynamic communication demands, including large public events, sudden traffic surges, or routine daily traffic variations [2, 25]. By adapting to these fluctuations, HAPS alleviate congestion, enhance quality of service (QoS), and improve energy efficiency. Compared to small cell densification, which requires extensive planning, fixed infrastructure, and offers limited flexibility, HAPS provide a faster and more cost-effective solution that can be deployed in real time without physical constraints. A QoS-driven deployment framework was proposed in [25] to optimize HAPS operation based on coverage, capacity, and latency metrics, balancing efficiency and cost. In integrated networks where ground and aerial BSs share spectrum, interference remains a key issue. As noted in [26], advanced techniques such as dynamic resource allocation, spectrum sharing, and adaptive power control enable efficient HAPS operation and seamless coexistence with TNs, ensuring robust performance under variable traffic conditions.
III-B Network Infrastructure and Management
III-B1 Backhauling
@@
Network backhaul is a fundamental component of communication infrastructure, responsible for high-capacity, low-latency data transmission between BSs and the core network. In dense urban areas, fiber optics remain the dominant solution, offering high throughput and minimal latency, whereas copper cables, limited in bandwidth, are largely obsolete for modern traffic demands. In contrast, rural and remote regions often rely on satellite or microwave links, as fiber deployment is economically and logistically prohibitive [27]. Non-geostationary satellite systems require costly ground gateways with precise placement to maintain reliable connectivity [28].
Although wireless backhauling has been studied extensively [29], HAPS are emerging as a promising enabler for next-generation networks. Positioned in the stratosphere, they provide near-LoS connectivity with lower latency compared to satellites, supporting applications such as small cell backhauling, wireless sensor networks, and Earth observation [30]. In-band mmWave backhaul has been demonstrated [31], while multi-band approaches for enhanced capacity are discussed in [32].
@@
Recent studies propose advanced architectures leveraging HAPS. In [33], a cell-free scheme introduces HAPS as an aerial central processing unit (CPU) aggregating UAVs via terahertz (THz) links, where large antenna arrays mitigate path loss. Hybrid radio frequency (RF)/FSO solutions further enhance flexibility: [34] presents a cooperative hybrid ARQ protocol ensuring latency fairness under varying conditions, and [35] explores hybrid LEO-to-HAPS FSO and HAPS-to-ground RF links. Collectively, these advancements demonstrate the potential of HAPS-enabled backhaul to deliver scalable, low-latency connectivity across heterogeneous environments.
III-B2 Handover Management
Frequent handovers and associated failures represent a major challenge in satellite communication systems, particularly for fast-moving non-geostationary satellites that require continuous connectivity updates. HAPS mitigate these issues in two key roles: (i) as intermediaries between satellites and UEs, handling handover requests on behalf of satellites, and (ii) as standalone communication managers, directly managing UE connections. In both cases, HAPS improve network stability and reduce signaling overhead.
In [36], leveraging HAPS as intermediaries was shown to decouple satellite mobility from ground-based handover procedures, simplifying management and enhancing continuity. Dynamic LEO-HAPS handover strategies proposed in [37] achieved lower delays, signaling costs, and drop rates than traditional 5G methods. Similarly, [30] explored FSO-based LEO-HAPS links, demonstrating high-speed, low-latency handovers even under high satellite mobility.
@@ III-B2 Handover Management
When operating as quasi-stationary communication nodes, HAPS significantly reduce handover frequency for terrestrial UEs, providing seamless service in high-mobility scenarios (e.g., vehicles, aircraft). Their stable positioning eliminates inter-platform handovers common in LEO constellations. The airliner system in [38], combining beamforming for HAPS-to-ground and FSO backhaul to LEO/ medium Earth orbit (MEO) satellites, demonstrated efficient handover management even in dynamic environments. Moreover, optical HAPS-satellite links, as analyzed in [39], further enhance reliability by ensuring robust backhaul connectivity, thereby minimizing disruptions in HAPS-to-UE sessions.
III-C IoT and Emerging Applications
III-C1 IoT sensing
@@
HAPS can significantly enhance IoT sensing and monitoring by providing wide, stable coverage and integrating edge computing for low-latency data processing close to the source. Their dynamic positioning supports optimized sensor deployment and efficient data acquisition, making them ideal for remote and critical environments.
@@
A reliable IoT architecture leveraging HAPS was proposed in [40], integrating narrow band IoT (NB-IoT) and long range (LoRa) to deliver scalable, cost-efficient connectivity across underserved regions. Operating at 17–35 km, HAPS achieved coverage up to 600 km, surpassing terrestrial solutions in scalability and redundancy through overlapping cells.
In [41], a joint HAPS-UAV framework for IoT applications—especially disaster scenarios—was introduced. Combining umbrella coverage from HAPS and agile UAVs, the system ensured continuous service with sub-millisecond latency (0.07–0.4 ms for HAPS), while NB-IoT/LoRa access schemes minimized collisions under dense deployments.
Complementary LoRa-based monitoring was examined in [42], which optimized parameters such as altitude, transmission power, and device density under shadowing and collision constraints, confirming the viability of HAPS for large-scale environmental IoT.
@@
To enhance HAPS network capacity, [43] introduced a 5G IoT architecture with a uniform rectangular array antenna and a Long Short-Term Memory (LSTM) model for direction-of-arrival prediction. The approach improved beamforming precision and link stability, indirectly benefiting IoT reliability. Joint LEO-HAPS IoT integration was explored in [44], where HAPS relayed data to ground centers via LEO satellites, optimizing throughput and power consumption under QoS and storage constraints. Similarly, [45] presented a smart rural framework combining 5G, cloud, IoT, and HAPS, enabling automated remote data access and processing.
III-C2 Mobile Edge Computing and Fast UE Movement
@@
MEC enhances network performance by placing computing resources near UEs, reducing latency for real-time services. However, maintaining seamless connectivity for fast-moving UEs requires advanced handover and predictive mechanisms to sustain service continuity. In [46], the integration of HAPS and UAVs within Vehicular Edge Computing (VEC) systems was analyzed using queuing theory to optimize task offloading in dynamic vehicular environments. The study showed that HAPS, with quasi-stationary operation, broad coverage, and strong LoS links, outperform UAVs by supporting larger computing payloads and offering lower latency, making them ideal for large-scale, high-mobility VEC/MEC scenarios. Although prior works such as [47, 48] optimized UAV-enabled MEC for task completion and energy efficiency, these approaches can be extended to HAPS, which offer superior payload capacity and operational stability. Yet, research on HAPS-based MEC remains limited, representing a promising avenue for future exploration.
III-C3 Aerial Data Centers and AR/VR Support
@@
Beyond edge computing, HAPS can function as aerial data centers, delivering centralized cloud services for data storage, processing, and distribution from the stratosphere. With wide coverage, high capacity, and low-latency connectivity, they support resource-intensive applications such as AR and VR [49], enabling immersive experiences even in remote or underserved regions [50]. In [51], a HAPS-enabled VEC system for rural areas was analyzed, showing performance gains through optimized trade-offs between computational load and latency. A cloud-integrated HAPS system proposed in [52] achieved up to 14% energy savings compared to terrestrial data centers, reducing offloading outages and delays in distributed server networks while supporting services such as IoT, vehicular networks, and social media. Furthermore, [44] introduced a collaborative LEO-HAPS framework in which HAPS acted as relays, forwarding IoT data to ground processing centers with optimized data rates, power efficiency, and QoS compliance.
III-D Advanced Control and Special Applications
III-D1 URLLC Support and Remote Driving Control
@@
Operating at high altitudes but closer to the Earth than traditional satellites, HAPS offer a unique combination of wide-area coverage, low-latency communication, and substantial payload capacity. These features enable the integration of advanced antenna systems and onboard processing units, making HAPS well-suited for supporting URLLC services such as remote driving control. In [53], a heterogeneous network integrating HAPS, terrestrial BSs, and autonomous vehicles was analyzed, showing that multi-connectivity (MC) is essential to meet stringent URLLC requirements. HAPS acted as auxiliary links sustaining data rates above 700 kbps with high reliability under dynamic conditions. Building on this, [54] extended the model with LEO satellites, optimizing MC path selection to satisfy QoS constraints with minimal link usage. Results revealed that LoS interference is a key performance-limiting factor in radio access links, which was further addressed in [55] through refined MC strategies for enhanced reliability and efficiency. Beyond URLLC, HAPS can improve Global Navigation Satellite System (GNSS) performance due to their quasi-stationary position and strong signal power. In [56], a HAPS-aided GPS architecture demonstrated improved 3D positioning accuracy, enhancing navigation and control functions. This capability is particularly valuable for remote or underserved regions—such as mountainous, coastal, or oceanic areas—where HAPS can act as super macro base stations (HAPS-SMBS) to backhaul isolated nodes and extend coverage [2].
III-D2 UAV Control and Aerial Highways
@@
A promising solution involves establishing aerial highways—structured UAV corridors that enable regulated, safe, and efficient airspace management [64, 65]. To support seamless communication along these corridors, recent works have optimized cellular network configurations and resource allocation strategies tailored for aerial traffic [66, 67, 68, 69, 70], while theoretical frameworks have been proposed to design structured airspace operations [71, 72].
@@
With their quasi-stationary position, large coverage footprint, and low-latency links, HAPS are well-suited to serve as backbone infrastructure for both open-air and corridor-based UAV networks. As discussed in Section III-B, HAPS significantly reduce handover frequency and ensure continuous connectivity for high-mobility UAVs, while avoiding the inter-platform coordination challenges inherent to LEO systems. This combination of coverage, stability, and operational efficiency positions HAPS as a key enabler for scalable and reliable UAV control.
III-E Energy and Sustainability

### v2

@@ IV HAPS Use Cases
The unique characteristics of HAPS enable a wide range of applications across telecommunications and beyond. In telecommunications, they primarily aim to extend coverage (Section IV-A) and enhance network efficiency (Section IV-B). Their rapid deployment and adaptability make them essential for bridging connectivity gaps in remote, underserved, or disaster-affected regions. Serving as mobile urban BSs or backhaul providers, HAPS improve capacity, reliability, and energy efficiency. Beyond communications, they support advanced applications (Section IV-C), including real-time IoT sensing, immersive augmented reality (AR)/ virtual reality (VR), precision control via URLLC (Section IV-D), and energy optimization (Section IV-E). By bringing processing closer to UEs, HAPS reduce latency and enable innovations across sectors such as agriculture, logistics, and disaster management. This section explores their roles in connectivity enhancement and advanced use cases shaping next-generation networks.
IV-A Connectivity Enhancement
IV-A1 Remote and Rural Area Coverage
@@
Collaboration between HAPS and LEO satellites was explored by Jia et al. [109], who formulated a mixed-integer nonlinear programming problem to enhance large-scale access and backhaul. The proposed restricted trilateral matching algorithm improved coordination, connectivity, and efficiency in remote regions. Zhao et al. [249] analyzed HAPS-LAPS integration using stochastic geometry, introducing a framework to assess platform density and altitude effects under backhaul constraints, offering design guidance for robust aerial networks in rural areas. Arum et al. [30] considered HAPS for extending rural coverage with low UE density through integration with TNs, emphasizing coexistence via dynamic resource management, interference mitigation, and techniques such as advanced antennas, radio environment maps, and device-to-device communication. Zhang, Kishk and Alouini [246] optimized resource allocation for heterogeneous services in rural HAPS networks using free space optical (FSO) backhaul. It proposed static and dynamic allocation schemes leveraging DRL, improving age of information, data rates, and overall network efficiency.
IV-A2 Natural Disaster Region Coverage
@@
In the aftermath of natural disasters, when TNs collapse, HAPS can provide real-time aerial surveillance and rapid, reliable communication. Covering vast and remote regions, they ensure effective disaster response and emergency connectivity for first responders [75]. Yu et al. [238] examined UAV-based HAPS for disaster scenarios, highlighting their rapid deployment, flexible coverage, and communication recovery, while identifying challenges such as backhaul limitations, platform constraints, and communication capacity that require targeted solutions. Real-time monitoring using HAPS was demonstrated by Baraniello et al. [37], showing persistent, high-resolution imaging from altitudes around 20 km, enabling stable, continuous coverage of affected areas. In the study provided by Lähdekorpi et al. [129], HAPS restored services in a disrupted universal mobile telecommunications system (UMTS) network, achieving 70% throughput recovery when terrestrial BSs were unavailable, underscoring their effectiveness in disaster recovery operations.
IV-A3 Unexpected Traffic Events
@@
The inherent mobility and on-demand deployment of HAPS make them well suited for managing dynamic communication demands, including large public events, sudden traffic surges, or routine daily traffic variations [116, 61]. By adapting to these fluctuations, HAPS alleviate congestion, enhance quality of service (QoS), and improve energy efficiency. Compared to small cell densification, which requires extensive planning, fixed infrastructure, and offers limited flexibility, HAPS provide a faster and more cost-effective solution that can be deployed in real time without physical constraints. A QoS-driven deployment framework was proposed by Dong et al. [61] to optimize HAPS operation based on coverage, capacity, and latency metrics, balancing efficiency and cost. In integrated networks where ground and aerial BSs share spectrum, interference remains a key issue. As noted by Wei et al. [227], advanced techniques such as dynamic resource allocation, spectrum sharing, and adaptive power control enable efficient HAPS operation and seamless coexistence with TNs, ensuring robust performance under variable traffic conditions. Jeon et al. [108] installed a reconfigurable intelligent surface (RIS) on a HAPS to provide extensive LoS and full-area coverage for UAV-BSs deployed in response to sudden urban traffic spikes. The combined optimization of RIS phases, array-partitioning algorithms, and positioning enhanced the energy efficiency of each UAV-BS.
IV-B Network Infrastructure and Management
IV-B1 Backhauling
@@
Network backhaul is a fundamental component of communication infrastructure, responsible for high-capacity, low-latency data transmission between BSs and the core network. In dense urban areas, fiber optics remain the dominant solution, offering high throughput and minimal latency, whereas copper cables, limited in bandwidth, are largely obsolete for modern traffic demands. In contrast, rural and remote regions often rely on satellite or microwave links, as fiber deployment is economically and logistically prohibitive [56]. However, non-geostationary satellite systems require costly ground gateways with precise placement to maintain reliable connectivity [34]. Although wireless backhauling has been studied extensively [174], HAPS are emerging as a promising enabler for next-generation networks. Positioned in the stratosphere, they provide near-LoS connectivity with lower latency compared to satellites, supporting applications such as small cell backhauling, wireless sensor networks, and Earth observation [158]. In-band mmWave backhaul has been demonstrated by Taori and Sridharan [207], while multi-band approaches for enhanced capacity were discussed by Tezergil and Onur [211].
@@
Recent studies proposed advanced architectures leveraging HAPS. Abbasi and Yanikomeroglu [13], introduced a cell-free scheme with HAPS as an aerial central processing unit (CPU) aggregating UAVs via terahertz (THz) links, where large antenna arrays mitigate path loss. Hybrid radio frequency (RF)/FSO solutions further enhanced flexibility: Dang, Le, Nguyen, and Pham [58] presented a cooperative hybrid ARQ protocol ensuring latency fairness under varying conditions, and Yahia, Erdogan, and Kurt [233] explored hybrid LEO-to-HAPS FSO and HAPS-to-ground RF links. Collectively, these advancements demonstrate the potential of HAPS-enabled backhaul to deliver scalable, low-latency connectivity across heterogeneous environments.
IV-B2 Handover Management
@@
Frequent handovers and associated failures represent a major challenge in satellite communication systems, particularly for fast-moving non-geostationary satellites that require continuous connectivity updates. HAPS mitigate these issues in two key roles: (i) as intermediaries between satellites and UEs, handling handover requests on behalf of satellites, and (ii) as standalone communication managers, directly managing UE connections. In both cases, HAPS improve network stability and reduce signaling overhead. Leveraging HAPS as intermediaries was shown by Cianca et al. [49] to decouple satellite mobility from ground-based handover procedures, simplifying management and enhancing continuity. Dynamic LEO-HAPS handover strategies proposed by Li et al. [131] achieved lower delays, signaling costs, and drop rates than traditional 5G methods. Prior to that, Mohammed, Mehmood, Pavlidou, and Mohorcic [158] explored FSO-based LEO-HAPS links, demonstrating high-speed, low-latency handovers even under high satellite mobility. When operating as quasi-stationary communication nodes, HAPS significantly reduce handover frequency for terrestrial UEs, providing seamless service in high-mobility scenarios (e.g., vehicles, aircraft). Their stable positioning eliminates inter-platform handovers common in LEO constellations. Combining beamforming for HAPS-to-ground and FSO backhaul to LEO/MEO satellites, Srinivasan et al. [202] demonstrated efficient handover management in the airliner system even in dynamic environments. Moreover, optical HAPS-satellite links, as analyzed be Perlot et al. [173], can further enhance reliability by ensuring robust backhaul connectivity, thereby minimizing disruptions in HAPS-to-UE sessions.
IV-C IoT and Emerging Applications
IV-C1 IoT sensing
@@
Regarding this, HAPS can significantly enhance IoT sensing and monitoring by providing wide, stable coverage and integrating edge computing for low-latency data processing close to the source. Their dynamic positioning supports optimized sensor deployment and efficient data acquisition, making them ideal for remote and critical environments.
@@
A reliable IoT architecture leveraging HAPS was proposed by Sibiya and Olugbara [196], integrating narrow band IoT (NB-IoT) and long range (LoRa) to deliver scalable, cost-efficient connectivity across underserved regions. Operating at 17–35 km, HAPS achieved coverage up to 600 km, surpassing terrestrial solutions in scalability and redundancy through overlapping cells.
@@
Andreadis, Giambene, and Zambon [27] introduced a joint HAPS-UAV framework for IoT applications—especially disaster scenarios. Combining umbrella coverage from HAPS and agile UAVs, the system ensured continuous service with sub-millisecond latency (0.07–0.4 ms for HAPS), while NB-IoT/LoRa access schemes minimized collisions under dense deployments. Complementary LoRa-based monitoring was examined by Giambene and Korre [76], which optimized parameters such as altitude, transmission power, and device density under shadowing and collision constraints, confirming the viability of HAPS for large-scale environmental IoT. To enhance HAPS network capacity, Xiao, Li, and Zhao [231] introduced a 5G IoT architecture with a uniform rectangular array antenna and a Long Short-Term Memory (LSTM) model for direction-of-arrival prediction. The approach improved beamforming precision and link stability, indirectly benefiting IoT reliability. Joint LEO-HAPS IoT integration was explored by Ei et al. in [63], where HAPS relayed data to ground centers via LEO satellites, optimizing throughput and power consumption under QoS and storage constraints. Similarly, Monzon and Alvarez [160] presented a smart rural framework combining 5G, cloud, IoT, and HAPS, enabling automated remote data access and processing.
IV-C2 Mobile Edge Computing and Fast UE Movement
@@
MEC enhances network performance by placing computing resources near UEs, reducing latency for real-time services. However, maintaining seamless connectivity for fast-moving UEs requires advanced handover and predictive mechanisms to sustain service continuity. Traspadini et al. [214] analyzed the integration of HAPS and UAVs within Vehicular Edge Computing (VEC) systems using queuing theory to optimize task offloading in dynamic vehicular environments. The authors showed that HAPS, with quasi-stationary operation, broad coverage, and strong LoS links, outperform UAVs by supporting larger computing payloads and offering lower latency, making them ideal for large-scale, high-mobility VEC/MEC scenarios. Later, Seid et al. [188] proposed to use aerial MEC servers to support ground user equipments and emergency communications. The authors introduced a hierarchical deep federated learning resource allocation (RA) scheme for intelligent aerial-enabled smart city cyber-physical system, aiming at optimizing energy consumption while preserving task offloading privacy and maintaining QoS in dynamic environments.Although prior works such as [204, 225] optimized UAV-enabled MEC for task completion and energy efficiency, these approaches can be extended to HAPS, which offer superior payload capacity and operational stability. Yet, research on HAPS-based MEC remains limited, representing a promising avenue for future exploration.
IV-C3 Aerial Data Centers and AR/VR Support
@@
Beyond edge computing, HAPS can function as aerial data centers, delivering centralized cloud services for data storage, processing, and distribution from the stratosphere. With wide coverage, high capacity, and low-latency connectivity, they support resource-intensive applications such as AR and VR [20], enabling immersive experiences even in remote or underserved regions [126]. Traspadini et al. [213] presented a HAPS-enabled VEC system for rural areas, showing performance gains through optimized trade-offs between computational load and latency. A cloud-integrated HAPS system proposed in [153] achieved up to 14% energy savings compared to terrestrial data centers, reducing offloading outages and delays in distributed server networks while supporting services such as IoT, vehicular networks, and social media. Furthermore, [63] introduced a collaborative LEO-HAPS framework in which HAPS acted as relays, forwarding IoT data to ground processing centers with optimized data rates, power efficiency, and QoS compliance.
IV-D Advanced Control and Special Applications
IV-D1 URLLC Support and Remote Driving Control
@@
Operating at high altitudes but closer to the Earth than traditional satellites, HAPS offer a unique combination of wide-area coverage, low-latency communication, and substantial payload capacity. These features enable the integration of advanced antenna systems and onboard processing units, making HAPS well-suited for supporting URLLC services such as remote driving control. Salehi et al. [187] analyzed a heterogeneous network integrating HAPS, terrestrial BSs, and autonomous vehicles, showing that multi-connectivity (MC) is essential to meet stringent URLLC requirements. HAPS acted as auxiliary links sustaining data rates above 700 kbps with high reliability under dynamic conditions. Building on this, the authors in [186] extended the model with LEO satellites, optimizing MC path selection to satisfy QoS constraints with minimal link usage. Results revealed that LoS interference is a key performance-limiting factor in radio access links, which was further addressed by Manzoor, Ozger, Schupke, and Cavdar [150] through refined MC strategies for enhanced reliability and efficiency. Beyond URLLC, HAPS can improve Global Navigation Satellite System (GNSS) performance due to their quasi-stationary position and strong signal power. Zheng et al. [250] demonstrated a HAPS-aided GPS architecture that improves 3D positioning accuracy, enhancing navigation and control functions. This capability is particularly valuable for remote or underserved regions—such as mountainous, coastal, or oceanic areas—where HAPS can act as super macro base stations (HAPS-SMBS) to backhaul isolated nodes and extend coverage [116].
IV-D2 UAV Control and Aerial Highways
@@
A promising solution involved establishing aerial highways—structured UAV corridors that enable regulated, safe, and efficient airspace management [44, 40]. To support seamless communication along these corridors, recent works have optimized cellular network configurations and resource allocation strategies tailored for aerial traffic [38, 39, 145, 47, 197], while theoretical frameworks have been proposed to design structured airspace operations [118, 117].
@@
With their quasi-stationary position, large coverage footprint, and low-latency links, HAPS are well-suited to serve as backbone infrastructure for both open-air and corridor-based UAV networks. As discussed in Section IV-B, HAPS significantly reduce handover frequency and ensure continuous connectivity for high-mobility UAVs, while avoiding the inter-platform coordination challenges inherent to LEO systems. This combination of coverage, stability, and operational efficiency positions HAPS as a key enabler for scalable and reliable UAV control.
IV-E Energy and Sustainability